Hello, im having difficulty with this problem.
How do you graph this linear equation
x+y=12
8x+6y=86
PLease help I keep getting decimals it doesn't seem right.
Thank You

Answer by lwsshak3(11628)

How do you graph this linear equation
x+y=12
8x+6y=86
..
6x+6y=72
8x+6y=86
subtract
-2x=-14
x=7
y=12-x
y=5
point of intersection: (7,5)
Both equations are straight lines that intersect at (7,5)